How many liters of gas can be collected at 25 C and 0.98atm if 0.063 grams of zinc metal fully reacts in a solution of hydrochloric acid.
1
Expert's answer
2020-04-14T01:43:13-0400

Zn+2HCl=ZnCl2+H2

0.063/65=0.000969230769 moles.

V=nRT/p= 0.000969230769 x 0.082 x 298 / 0.98 = 0.0241674726 L.